OpenStudy (eliesaab):

What is the largest positive integer \(n\) such that \(n^3+100\) is divisible by \( n+10\)?

OpenStudy (solomonzelman):

n=2 ? Trial and error.

OpenStudy (solomonzelman):

(Just noticed that 12 divides 108)

OpenStudy (solomonzelman):

Oh sorry, the largest:)

OpenStudy (seratul):

Wouldn't there be an endless amount of answers?

OpenStudy (eliesaab):

No. There is on;y one answer

OpenStudy (solomonzelman):

\(\displaystyle \frac{ n^3+100}{n+10}=\frac{ n^3+1000-900}{n+10}=(n^2-10n+100)-\frac{900}{n+10}\)

OpenStudy (solomonzelman):

so, we want the largest integer \(n\) such that \(900\) divides \(n+10\). \(\color{red}{n=890}\).

OpenStudy (eliesaab):

Excellent method

OpenStudy (solomonzelman):

TY:)

OpenStudy (solomonzelman):

Nice problem, I liked it:)